Make New Entry,
Make Followup Entry
User name A. Camsonne
Log entry time 03:03:27 on February 9,2009
Entry number 259912
keyword=Compton rates seems OK, but odd behavior of GSO
No signal could be seen when beam came back, even on the scope. So was not sure if the detector was flooded with background. DAQ was checked out to work ok with pulser trigger and triggering on noise.
After asking to MCC to improve beam profile, it seems it improved a bit since signal could be seen but not sure if this related. I could see signal correlated with cavity by increasing gain and HV. So for now rates looks ok and some signal could be seen, I might try a vertical scan to improve the signal. Will double check the cabling and signal of the detector at next access too.